OpenStudy (ammarah):

A polynomial f and a factor of f are given. Factor f completely. f(x)=x^3-3x^2-16x-12;x-6

ganeshie8 (ganeshie8):

familiar wid synthetic division ?

OpenStudy (ammarah):

yes i got (x-6)(x+2)(x+1) is that correct?

ganeshie8 (ganeshie8):

Yes ! good job :)

OpenStudy (ammarah):

Im having trouble with this problem f(x)=4x^4+26x^3-8x^2+39x-21;x+7

ganeshie8 (ganeshie8):

use synthetic division, divide x+7

OpenStudy (ammarah):

i am but its turning out to be big numbers and not equal to 0

ganeshie8 (ganeshie8):

-7 | 4 26 -8 39 -21 | -------------------------------------- 4

ganeshie8 (ganeshie8):

lets work it and see wat we get

OpenStudy (ammarah):

-2, -22,238

ganeshie8 (ganeshie8):

-7 | 4 26 -8 39 -21 | -28 -------------------------------------- 4 -2

ganeshie8 (ganeshie8):

fine so far ?

ganeshie8 (ganeshie8):

-7 | 4 26 -8 39 -21 | -28 14 -42 21 -------------------------------------- 4 -2 6 -3 | 0

OpenStudy (ammarah):

ohh wait nvm i calculated it wrong

ganeshie8 (ganeshie8):

so it will be :- (x+7) (4x^3 - 2x^2 + 6x - 3)

ganeshie8 (ganeshie8):

:)

OpenStudy (ammarah):

(x+7)(4x^3-2x^2+6x-3)

OpenStudy (ammarah):

yes

OpenStudy (ammarah):

so do i have to factor it completely....like how

ganeshie8 (ganeshie8):

yes you will have to factor it completely

ganeshie8 (ganeshie8):

familiar wid rational root theorem ?

OpenStudy (ammarah):

yeah but i kinda forgot...is it the one where u list the numbers

ganeshie8 (ganeshie8):

4x^3-2x^2+6x-3

ganeshie8 (ganeshie8):

yes, ``` rational roots will be of form p/q p = factors of constant term (3 here) q = factors of leading coeffecient (4 here) ```

ganeshie8 (ganeshie8):

factors of 3 = ? factors of 4 = ?

OpenStudy (ammarah):

1,3 1,2,2,4

OpenStudy (ammarah):

plus and minus

ganeshie8 (ganeshie8):

yes, so possible rational roots are :- \(\pm( 1, ~3, ~2, ~4, ~1/2, ~1/4, ~3/2,~ 3/4)\)

OpenStudy (ammarah):

ok...then what

ganeshie8 (ganeshie8):

start by testing all possible POSITIVE roots first start wid +1

OpenStudy (ammarah):

ohhh

ganeshie8 (ganeshie8):

4x^3-2x^2+6x-3 put x = 1 above, do u get 0 ?

OpenStudy (ammarah):

wait do use the 4,-2,6,-3?

ganeshie8 (ganeshie8):

dint get u :|

OpenStudy (ammarah):

like do i use the synthetic division with 4 -2 6 -3 | 0 or with the original which is 4,26,-8,39,-21

ganeshie8 (ganeshie8):

we got the depressed cubic equation, so we need to work on this cubic oly use 4 -2 6 -3

OpenStudy (ammarah):

ok so 1 doesnt work

ganeshie8 (ganeshie8):

wat about 2, 3, 4 ?

ganeshie8 (ganeshie8):

does any of it work ? i dont think so... try \(1/2\)

OpenStudy (ammarah):

yeah it works

ganeshie8 (ganeshie8):

good, whats the depressed quadratic after division wid -1/2 ?

OpenStudy (ammarah):

4x+6

ganeshie8 (ganeshie8):

no, you should get a quadratic. show me ur synthetic division

OpenStudy (ammarah):

4x^2+0+6x

ganeshie8 (ganeshie8):

oh you mean 4x^2 + 6 ?

OpenStudy (ammarah):

yeah yeah im sorry

ganeshie8 (ganeshie8):

its 4x^2 + 0x + 6 ok

ganeshie8 (ganeshie8):

np :) so, 4x^3-2x^2+6x-3 = (x-1/2)(4x^2+6)

OpenStudy (ammarah):

yes so its (x+7)(x-1/2)(4x^2+6)

ganeshie8 (ganeshie8):

4x^4+26x^3-8x^2+39x-21 = (x+7)(x-1)(4x^2 + 6)

ganeshie8 (ganeshie8):

Yes ! dont leave it like that, u can factor further the last quadratic factor aswell

ganeshie8 (ganeshie8):

4x^2 + 6 = 2(2x^2+3) = ?

OpenStudy (ammarah):

idk..

OpenStudy (ammarah):

how would u factor that

ganeshie8 (ganeshie8):

use quadratic formula and find zeroes of 2x^2 + 3

ganeshie8 (ganeshie8):

a = 2 b = 0 c = 3

ganeshie8 (ganeshie8):

\(\large x = \frac{-b \pm \sqrt{b^2-4ac}}{2a}\)

OpenStudy (ammarah):

i got